a recursive function is defined by can=2an-1 with a0=1. find the value of a3


1
Expert's answer
2021-09-07T18:01:21-0400
"a_0=1"

"ca_1=2a_0=>a_1=\dfrac{2}{c}"

"ca_2=2a_1=>a_2=\dfrac{4}{c^2}"

"ca_3=2a_2=>a_3=\dfrac{8}{c^3}"